The gravitational force is thrice as great on a 3 kg rock as on a 1 kg rock. What should you do so that the 3 kg will not fall with thrice the acceleration as that of the 1 kg rock?


1
Expert's answer
2021-03-21T11:26:54-0400

If the gravitational force is directly proportional to the mass, the aceleration should be inversly proportional to the mass, in order the 3 kg will not fall with thrice the acceleration as that of the 1 kg rock. And it is true, since the second Newton's law holds:


"a = \dfrac{F}{m}"

Let "F_1, m_1 = 1kg" be the gravitational force and mass of the first rock respectively. Then "F_2 = 3F_1, m_2 = 3m_1" are the gravitational force and mass of the second rock respectively. Then the accleleration of the first rock is:


"a_1 = \dfrac{F_1}{m_1}"

And the acceleration of the secon force is:


"a_2 = \dfrac{F_2}{m_2} = \dfrac{3F_1}{3m_1} = \dfrac{F_1}{m_1} = a_1"

Thus, both accelerations are equal.


Answer. The aceleration should be inversly proportional to the mass.